Grade 12Algebra

The projections of a vector on the three coordinate axis are 6, – 3, 2 respectively. The direction cosines of the vector are -

saket kumar
12 Years agoGrade 12

1 Answer

SHAIK AASIF AHAMED
12 Years ago
the directional cosines of the vector are given by\frac{6}{\sqrt{6^{2}+3^{2}+2^{2}}},\frac{-3}{\sqrt{6^{2}+3^{2}+2^{2}}},\frac{2}{\sqrt{6^{2}+3^{2}+2^{2}}},
\frac{6}{7},\frac{-3}{7},\frac{2}{7}
Thanks and Regards
Shaik Aasif
askIITians Faculty